MONT 102N – Modeling the Environment Debating the Issues November 21, 2011

General Information

As announced in the course syllabus and the semester schedule on the course homepage, during the final week of the semester (class meetings of December 5, 7, and 9), we will be stepping back from the mathematics we have studied to reconsider some larger environmental issues we have touched on. We will do this with a series of “debates” based on recent opinion pieces published in the New York Times series Room For Debate. Each of the links below leads to a group of related statements by people with various opinions about the overall question.

The Assignment

For each class, the assigned group of students will begin by presenting the views of the various people who wrote the Room For Debate pieces (see below for your assignment). Note: Your job is to try to understand your assigned writer’s point of view and explain it clearly and succinctly (even if you do not agree with it yourself!) Please don’t just read the writer’s piece to the class. The idea is to present what they are saying after having thought it through yourself and having anticipated possible responses, etc.

You will want to consult with the other students in your group so that you all understand the key issues, you all have a good handle on what each of the writers is saying, and you see how the opinions differ from each other and connect with each other. Plan on taking about 5 minutes for laying out each of these viewpoints (this will be roughly 25 to 30 minutes in all). Then, the floor will be open for questions from Prof. Little and the other students in the class. Your job during this phase is to try to formulate answers to questions as you think your assigned writer would. To be really well-prepared for questions and understand some of the issues, you may want to do some background reading to have other facts at your disposal that support your writer’s points. At the end of the period, I will give everyone a chance to say what they thought about their assigned writer’s point of view. How hard or easy was it to argue this way? Did this experience change your mind at all (or perhaps give you new insight into a position different from your own)?

Sources

Your main source will be the series of short opinion pieces from the New York Times web site, found at the links given below. The College has an institutional electronic subscription to the New York Times web site which should allow you unlimited access to these and other articles from computers in any one of the college libraries. Consult with our Science Librarian, Ms. Barbara Merolli if you need help accessing the materials for your topic.

Important Note

For typographical reasons, the http://www.nytimes.com/roomfordebate at the start of all three links has been omitted – you will need to include that, of course, when you look up your sources.

Topics

A. Can the planet support 10 billion people? This Room for Debate appeared in May 2011, in advance of the recent passing of the 7 billion mark by the Earth’s human population. Here is the link to the New York Times web site where the original “debate” was published:

/2011/05/04/can-the-planet-support-10-billion-people/

1. Luke B. – Jamais Cascio, How 10 Billion Can Survive

2. Andrew B. – Jason Clay, More Efficient Food Production

3. Katie B. – Joel Cohen, We Can Change the Future

4. Lyndsay D. – David Bloom, Africa’s Daunting Challenges

5. Mike D. – Warren Sanderson, Educate the Masses

6. Amanda F. – Brad Allenby, Technology’s Impact on Population

B. What if Republicans closed the EPA? Some of the Republican candidates for President in the upcoming 2012 election (though surprisingly not Rick Perry in his now-infamous “brain freeze” moment) have made no bones of the fact that the Environmental Protection Agency is on their list of Federal agencies targeted for closure. But what would the impact of that choice be? Here is the link to the series of opinion pieces you will be looking at:

/2011/08/24/what-if-republicans-closed-the-epa/

1. Matt G. – Jonathan Adler, Fixing, Not Ending Regulation

2. Abbey L. – Chip Jacobs, No E.P.A.? Welcome Back Smog

3. Colby L. – Peter Lehner, A Cop on the Beat

4. Fabiana M. – Marlo Lewis, Bureaucrats Have Gone Rogue

5. Drew M. – Rick Schenkman, Dismantling an Agency Isn’t Easy

C. What are the risks of the Keystone XL pipeline project? The Obama administration has recently delayed making a decision about approval for a large and environmentally sensitive pipeline that would carry crude oil from a newly-developed oil sands area in Alberta, Canada, to the U.S. Gulf Coast for refining and distribution. What are the risks and issues behind developing this resource? Here is the link for this topic:

/2011/10/03/what-are-the-risks-of-the-keystone-xl-pipeline-project

1. Abby M. – Lisa Hymas, Nebraskans vs. The State Department

2. Megan N. – James Goeke, The Truth About Aquifers

3. Maria P. – Adam Brandt, What Choice Do We Have?

4. Ed S. – Tim Bradner, Lessons From Alaska

5. Julie V. – Najmedin Meshkati, The Odds of Human Error

6. Jake Y. – Mohammad Najafi, Why Take The Risk?